Sn Capaety Osage Alter Bridge Mfg., Inc., is currently operating at only 94 percent of fixed asset capacity. Current sales are $540,000. Fixed assets are $420,000 and sales are projected to grow to $760,000. How much in new fixed assets are required to support this growth in sales? Assume the company wants to operate at full capacity. Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g. 32.16)) New fixed assets
Answer #1

Full Capacity Sales = 540000 / .94

Full Capacity Sales =$574468.085

Current percentage of fixed assets to full capacity sales: 420000 / 574468.085 = .73

Total dollar amount of fixed assets needed for sales: .73 * 760000 = $554800

New Fixed Assets = Total fixed assets at the new sales - Current level of fixed assets

New Fixed Assets = 554800 - 420000

New Fixed Assets = $134800